OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (-1, -3) and (2, 1)

OpenStudy (anonymous):

4/3 is slope

OpenStudy (anonymous):

try to find the slope using the formula \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

8/3+b=1

OpenStudy (anonymous):

b=?

OpenStudy (anonymous):



OpenStudy (anonymous):

b=-5/3

OpenStudy (anonymous):

y=4/3x-5/3

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@yguo17 do not give people answers directly... give them hints but dont write the answer

OpenStudy (anonymous):

we are here to help people learn how to do stuff so that when they are offline they can use the help to solve real life situations (classwork, work...)

OpenStudy (anonymous):

Or you can just find the slope using \[ m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] then use the point slope form, do some algebra and transform it into y=mx+b form \[y-y _{1}=m(x-x _{1})\] Where \[(x_1,y_1)\] are your first coordinates, and slope is m.